OpenStudy (anonymous):

Find the slope and the y-intercept of the line: 4x=9y + 63. The slope is? and the y-intercept is (0,?).

OpenStudy (aroub):

-9y=-4x+63 y=4/9x-7 slope= 4/9 y-intercept=-7

OpenStudy (anonymous):

(0,-7)?

OpenStudy (aroub):

exactly! :)

OpenStudy (anonymous):

You ROCK!